A certain metal is 40% copper. How many kilograms of this metal must be mixed with 80kg of a metal that is 70% copper to get a metal that is 50% copper?
Answer by stanbon(75887)

A certain metal is 40% copper. How many kilograms of this metal must be mixed with 80kg of a metal that is 70% copper to get a metal that is 50% copper?
-------------------
Equation:
copper + copper = copper
0.40x + 0.70*80 = 0.50(x+80)
-----
40x + 70*80 = 50x + 50*80
-10x = -20*80
----
x = 2*80
x = 160 kg (amt of 40% alloy needed)
=========================================
Cheers,
Stan H.
-----------------